OpenStudy (anonymous):

The temperature of a city changed by −15 degrees Celsius over a 5-week period. The temperature changed by the same amount each week. Part A: Write an expression to show the average change in the temperature of the city each week. Part B: Simplify the expression and explain, using words, what your answer means.

OpenStudy (anonymous):

@Michele_Laino @texaschic101

OpenStudy (michele_laino):

here, by definition, the average change is: \[\frac{{\Delta \theta }}{{\Delta t}} = \frac{{ - 15}}{5} = ...?\]

OpenStudy (michele_laino):

correct!

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

the meaning of your answer is given by the subsequent statement: "we can suppose an average change of temperature about -3 degrees for each week"

OpenStudy (anonymous):

so i write ^ @Michele_Laino

OpenStudy (michele_laino):

the simplification is: \[\frac{{ - 15}}{5} = - 3\]